OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 7 inches and its preimeter is 42 inches, find the length of the rectangle. The length of the rectangle is __ inches I know the formula is P=2L+ 2W but the length is needed do I use the same formula?

OpenStudy (anonymous):

P = 2L + 2W 42 = 2L + 2(7) 42 = 2L + 14 42 - 14 = 2L 28 = 2L 28/2 = L 14 = L check.... P = 2L + 2W 42 = 2(14) + 2(7) 42 = 28 + 14 42 = 42 (correct)

OpenStudy (mathstudent55):

You have perimeter = P = 42 in. width = W = 7 in. The only unknown is the length, L Use the formula you have, and replace P and W with the known numbers. Then solve for L, the length.